# Training objective drives the consistency of representational similarity across datasets

### v1

@@ Abstract
The Platonic Representation Hypothesis claims that recent foundation models are converging to a shared representation space as a function of their downstream task performance, irrespective of the objectives and data modalities used to train these models [35]. Representational similarity is generally measured for individual datasets and is not necessarily consistent across datasets. Thus, one may wonder whether this convergence of model representations is confounded by the datasets commonly used in machine learning. Here, we propose a systematic way to measure how representational similarity between models varies with the set of stimuli used to construct the representations. We find that the objective function is the most crucial factor in determining the consistency of representational similarities across datasets. Specifically, self-supervised vision models learn representations whose relative pairwise similarities generalize better from one dataset to another compared to those of image classification or image-text models. Moreover, the correspondence between representational similarities and the models’ task behavior is dataset-dependent, being most strongly pronounced for single-domain datasets. Our work provides a framework for systematically measuring similarities of model representations across datasets and linking those similarities to differences in task behavior.
@@ 1 Introduction
Representation learning has seen remarkable progress in recent years, with state-of-the-art models achieving or even surpassing human-level performance in a wide range of computer vision tasks [65, 18, 64, 85, 61]. The claim of The Platonic Representation Hypothesis is that with this progress in task performance, model representation spaces tend to approach each other irrespective of the models’ training data or architecture [35]. Even between models trained on different modalities, better downstream task performances appear to result in more similar representations. Two questions naturally arise from this phenomenon: Do pairwise similarities of model representations transfer from one dataset to another, and are they generally linked to the models’ downstream behavior?
@@
Investigating the correspondence between representation and behavior has a long history in representation learning and adjacent fields [e.g. 32, 74]. Two models can have very different intermediate layer representations even though they show the same task behavior [32, 60, 48]. However, if two models show different behavior, one can be certain that the output layer representations are different, because identical output representations map to the same behavior and, vice versa, different output layer representations map to different behavior. What remains unclear is how behavior (e.g., downstream task accuracy) is affected by the convergence of model representations (i.e., the representations are similar according to a chosen similarity measure). Does this relationship depend on the specific similarity measure at hand, or may it be determined by the nature of a dataset?
@@
The field lacks consensus on how to define (pairwise) representational similarity [cf. 45, 42, 81, 22, 74] and how to systematically measure its relationship to behavior [cf. 26, 60, 74, 48]. Here, we use similarity measures that are widely used in the community and have been empirically proven to be useful for measuring representational similarity of neural networks, such as CKA [42, 67] or RSA [45, 44]. However, pairwise representational similarity is usually measured for a single dataset and not across datasets. It may be the case that the representations of two vision models are highly similar for a dataset of raccoons but very different for a dataset of sunflowers — due to their different domains. Similarly, the representations of two models may exhibit a high similarity value for satellite data but can show a low similarity for images of fruit — because of the images’ different structure.
@@
Thus, it appears crucial to scrutinize the factors that determine pairwise representational similarity across sets of stimuli. Knowing about those factors will enable us to generalize pairwise representational similarity across datasets and more effectively find model sets that have learned a similar representation of the world. For a large set of diverse vision models, we examine their pairwise representational similarities for both structured and unstructured datasets from different domains, using the publicly available Visual Task Adaptation Benchmark [VTAB; 87]. Our contributions and findings are as follows:
@@
  • We find that the objective function is the most crucial factor for determining the consistency of pairwise representational similarities across datasets, whereas architecture, model size, and training data do not appear to be qualitatively meaningful factors.
  • Pairwise similarities in SSL models generalize more reliably across stimulus sets compared to image-text and supervised models, which show high variance in their consistency due to dataset dependence.
  • We demonstrate that our findings hold irrespective of whether the kernel in the similarity measure emphasizes local or global representational structure.
  • The pairwise similarities of model representations show a strong correlation with the models’ differences in task performance for single-domain datasets, while multi-domain datasets show highly variable and specialized datasets consistently low correlations.
@@ 2 Related work
Measuring the similarity of deep neural networks. Representational similarity measures aim to quantify to what extent the representations of two models corresponding to the same input overlap. In the field of machine learning, they have been used as a tool to understand deep neural networks [51, 63, 56], investigate human-machine alignment [74, 82], and as an objective for model distillation [79, 70, 88]. Various ways of measuring representational similarity have been proposed, including Canonical Correlation Analysis [CCA; 34, 66, 58], Singular Vector Canonical Correlation Analysis [SVCCA; 66], Centered Kernel Alignment [CKA; 42, 14], and Representational Similarity Analysis [RSA; 45], among others [e.g. 81, 20, 5, 16]. They can be distinguished from functional similarity measures, such as performance difference [39], error consistency [25], and true-positive agreement [29], which are based on model outputs rather than internal representations [39], or model stitching [4, 15, 57], which measures compatibility, not similarity, of internal representations [33].
@@
Importantly, similarity measures may differ in the invariances they evoke and the properties of representational spaces they capture [39]. For example, local structure may be better captured by measures based on nearest neighbors [35] or localized kernels [42], while other kernels may better capture global similarity. Some argue that similarity measures have to satisfy certain conditions, such as the triangle inequality and symmetry to qualify as measures of similarity [e.g. 81, 23, 30, 37], while others see value in relaxing the conditions of similarity [e.g. 9, 10, 74, 48]. Here, we use CKA, as it allows probing both local and global representation structures by varying the kernel, does not require the compared representations to be of the same dimension, and is widely used in the research community [e.g. 68, 55, 70, 88, 21, 6].
Drivers of similarity in vision models. In contrast to language processing systems [e.g. 8, 1, 76], computer vision models are trained with a wide range of different learning objectives. These include (a) supervised learning objectives using classification datasets such as ImageNet-1k [46], or ImageNet-21 [69]; (b) SSL objectives with a reconstruction or multi-view matching loss [12, 84, 11, 31]; and (c) objectives for jointly learning image and text representations [65, 36, 75, 86]. This makes them interesting candidate models for analyzing representational similarities.
@@
The trend to larger model sizes [e.g. 18, 1] combined with increased dataset sizes [e.g. 72, 64] and multi-modality [e.g. 65, 75, 86] appears to be leading to a convergence of model representations [35]. Similarity between otherwise identical networks trained from different random initializations increases with width [59, 42], as predicted by theoretical studies of properties of neural networks in the infinite width limit [50, 17, 83]. Huh et al. [35] hypothesize that convergence is further driven by task generality and inductive bias. Other works find that architecture and dataset [68] play a central role in determining the representation structure of a (pre-)trained model.
@@
Consistency of similarity. In addition to the similarity function itself, any comparison of representations depends on the layer from which the representations have been extracted [42, 68, 7] and the dataset for which the representations have been computed [16, 38, 7]. For example, measuring similarity for a subset of a data distribution can yield different similarities than measuring it for the whole dataset, albeit maintaining the ranking of similarities between models [7]. From a behavioral perspective, it has been shown that the ranking of the agreement of ImageNet classifiers pairs generalizes to out-of-distribution data [3]. Due to the limited scope of prior work, such as the inclusion of only a small number of models [38, 7], the driving factors of the consistency of representational similarities across vision datasets have not yet been identified.
@@ 3 Methods
Figure 1: Pairwise similarity analysis framework. Let ${\bm{A}}$ and ${\bm{B}}$ be two sets of stimuli and $\Theta$ and $\Phi$ be two sets of models. For each dataset, we extract representations from all functions within the two model sets. Subsequently, for each pair of models $(f_{\theta},f_{\phi})$, where $f_{\theta}\in\Theta$ and $f_{\phi}\in\Phi$, we compute the CKA similarities between their representations, yielding a similarity vector ${\bm{s}}_{\left({\bm{A}},\Theta,\Phi\right)}$. This vector can be displayed as a matrix, where each entry represents the similarity between two models for a single dataset. In the scatter plot, we contrast two such vectors computed on the same models but evaluated on different datasets. The Pearson correlation coefficient $\rho$ between the similarities quantifies the consistency of similarities across the two datasets. The distribution of $\rho$ across all dataset pairs indicates the stability of (relative) representational similarities across stimuli.
In this section, we propose a novel approach for analyzing the consistency of representational similarities across multiple datasets. Fig. 1 provides a schematic overview of this framework.
@@
Computing representational similarities for a dataset. We compute the similarity between two models $f_{\theta}$ and $f_{\phi}$ for a dataset by applying CKA [14, 42] to their vector representations ${\bm{Z}}^{f_{\theta}}$ and ${\bm{Z}}^{f_{\phi}}$. CKA computes the similarity based on the normalized Hilbert-Schmidt Independence Criterion [28], applied to the kernel matrices of both representations. Using CKA with a linear kernel focuses on global similarity structure, while an RBF kernel with small $\sigma$ measures local similarity structure [42, 2]. We compared their behavior (shown in Fig. 3 and in Appx. E) and observed similar trends for both kernels. Therefore, we use CKA with a linear kernel for the remainder of this paper if not mentioned otherwise.
@@
Computing the similarity requires the full kernel matrix over the dataset, which scales quadratically with the number of data points. Hence, the exact computation becomes intractable for large datasets. However, we demonstrate in Appx. C that a diverse subset of $10{,}000$ samples is sufficient to accurately estimate the similarities of two models for an ImageNet-scale dataset.
@@
Representational similarities across datasets. Because image representations depend on both a model and a dataset, the representational similarities between two models may vary across different datasets. To investigate this, we analyze the consistency of pairwise model similarities across multiple datasets. We compute the correlation coefficient for pairwise model similarities (measured for a large set of model pairs) between two datasets (i.e., can the pairwise model similarities for one dataset be expressed as a function of the similarities for a different dataset). In addition, we partition the models into specific sets determined by the models’ training factors.
@@
Specifically, we scrutinize how the pairwise similarities of models from two model sets, $\Theta$ and $\Phi$, correspond between two datasets, ${\bm{A}}$ and ${\bm{B}}$. By computing the pairwise representational similarities between all model pairs $(f_{\theta},f_{\phi})$, separately for each dataset, where $f_{\theta}\in\Theta$, $f_{\phi}\in\Phi$ and $f_{\theta}\neq f_{\phi}$, we obtain a similarity vector ${\bm{s}}\in\mathbb{R}^{k}$, where $k$ denotes the number of distinct model combinations. A similarity vector for two sets of models and a dataset ${\bm{X}}$ is then defined as follows:
${\bm{s}}_{\left({\bm{X}},\Theta,\Phi\right)}=\left(\mathrm{CKA}({\bm{Z}}^{f_{\theta}}_{\bm{X}},{\bm{Z}}^{f_{\phi}}_{\bm{X}})\mid f_{\theta}\in\Theta,f_{\phi}\in\Phi,f_{\theta}\neq f_{\phi}\right)^{\top}\in\mathbb{R}^{k}.$
@@
To quantify the consistency of similarities between two datasets ${\bm{A}}$ and ${\bm{B}}$, we use the Pearson correlation coefficient between the similarity vectors ${\bm{s}}_{({\bm{A}},\Theta,\Phi)}$ and ${\bm{s}}_{({\bm{B}},\Theta,\Phi)}$, i.e., $\rho\left({\bm{s}}_{({\bm{A}},\Theta,\Phi)},{\bm{s}}_{({\bm{B}},\Theta,\Phi)}\right)$. The Pearson correlation measures the degree to which the similarity trends between models are preserved across datasets, focusing on the relative positioning of model pairs rather than the absolute similarity values. This correlation measures how well the models’ representational similarities are preserved across datasets. A high Pearson correlation indicates that the ordering of model similarities of two model pairs remains stable between datasets. In other words, if two models $f_{\phi}$ and $f_{\theta}$ are more similar than $f_{\phi}$ and $f_{\psi}$ on one dataset, a high Pearson correlation coefficient indicates that this relationship transfers to the other dataset. Conversely, a low correlation value suggests that the models’ relationships shift significantly depending on the dataset, showing variability in processing different data distributions. Thus, the Pearson correlation coefficient is a meaningful indicator of how dataset-dependent or dataset-invariant the model similarities are.
@@
To assess the consistency of these correlations across dataset pairs, we examine the distribution of correlation coefficients for all possible dataset pairs within the set of available datasets $\mathcal{D}$:
${\bm{R}}(\Theta,\Phi)=\left\{\rho\left({\bm{s}}_{({\bm{A}},\Theta,\Phi)},{\bm{s}}_{({\bm{B}},\Theta,\Phi)}\right)\mid{\bm{A}},{\bm{B}}\in\mathcal{D},{\bm{A}}\neq{\bm{B}}\right\}.$
The distribution of the correlation coefficients across all dataset pairs indicates the consistency of the pairwise model similarities.

### v2

# Objective drives the consistency of representational similarity across datasets

@@ Abstract
The Platonic Representation Hypothesis claims that recent foundation models are converging to a shared representation space as a function of their downstream task performance, irrespective of the objectives and data modalities used to train these models (Huh et al., 2024). Representational similarity is generally measured for individual datasets and is not necessarily consistent across datasets. Thus, one may wonder whether this convergence of model representations is confounded by the datasets commonly used in machine learning. Here, we propose a systematic way to measure how representational similarity between models varies with the set of stimuli used to construct the representations. We find that the objective function is a crucial factor in determining the consistency of representational similarities across datasets. Specifically, self-supervised vision models learn representations whose relative pairwise similarities generalize better from one dataset to another compared to those of image classification or image-text models. Moreover, the correspondence between representational similarities and the models’ task behavior is dataset-dependent, being most strongly pronounced for single-domain datasets. Our work provides a framework for analyzing similarities of model representations across datasets and linking those similarities to differences in task behavior.
Machine Learning, ICML
@@ 1 Introduction
Representation learning has seen remarkable progress in recent years, with state-of-the-art models achieving or even surpassing human-level performance in a wide range of computer vision tasks (Radford et al., 2021; Dehghani et al., 2023; Oquab et al., 2024; Zhai et al., 2023; Muttenthaler et al., 2024). With this progress in task performance, model representation spaces tend to approach each other irrespective of the models’ training data or architecture (Huh et al., 2024). Even between models trained on different modalities, better task performances result in more similar representations. Two questions naturally arise from this phenomenon: Do pairwise similarities of model representations transfer from one dataset to another, and are they generally linked to the models’ downstream behavior?
@@
Investigating the correspondence between representation and behavior has a long history in representation learning and adjacent fields (e.g. Hermann & Lampinen, 2020; Sucholutsky et al., 2023). Two models can have different intermediate layer representations even though they show the same task behavior (Muttenthaler et al., 2023; Lampinen et al., 2024). However, if two models show different behavior, one can be certain that the output layer representations are different, since identical output representations map to the same behavior. What remains unclear is how behavior (e.g., downstream task accuracy) is affected by the convergence of model representations (i.e., the similarity of representations). Does this relationship depend on the similarity measure at hand, or may it be due to the nature of a dataset?
The field lacks consensus on how to define (pairwise) representational similarity (cf. Kriegeskorte et al., 2008; Kornblith et al., 2019a; Williams et al., 2021; Duong et al., 2023; Sucholutsky et al., 2023) and how to systematically measure its relationship to behavior (cf. Geirhos et al., 2021; Muttenthaler et al., 2023; Sucholutsky et al., 2023; Lampinen et al., 2024). Here, we use similarity measures that are widely used and have been empirically proven to be useful for measuring representational similarity of neural networks, such as Centered Kernel Alignment (CKA; Kornblith et al., 2019a; Raghu et al., 2021), and Representational Similarity Analysis (RSA; Kriegeskorte et al., 2008).
@@
However, pairwise representational similarity is usually measured for a single dataset and not across datasets. It may be that the representations of two vision models are highly similar for a dataset of raccoons but dissimilar for a dataset of sunflowers — due to their different domains. Similarly, the representations of two models may exhibit high similarity for satellite data but show low similarity for images of fruit — because of the images’ different structure.
@@
Thus, it appears crucial to scrutinize the factors that determine pairwise representational similarity across sets of stimuli. Knowing about those factors will enable us to generalize pairwise representational similarity across datasets and more effectively find model sets that have learned a similar representation of the world. For a large set of diverse vision models, we examine their pairwise representational similarities for both structured and unstructured datasets from different domains. Our contributions and findings are as follows:
@@
  • We find that the objective function is a crucial factor for determining the consistency of pairwise representational similarities across datasets, whereas architecture and model size do not appear to be major factors. In particular, SSL models show more reliable generalization across stimulus sets compared to image-text and supervised models.
  • Our findings show the same trend across different similarity measures, including those that emphasize local versus global representational structure.
  • Pairwise similarities of model representations strongly correlate with the models’ differences in task performance for single-domain datasets, while multi-domain datasets show highly variable and specialized datasets consistently low correlations.
@@ 2 Related work
Measuring the similarity of deep neural networks. Representational similarity measures quantify the degree of overlap between the representations of two models when processing the same input. They have been used as a tool to understand deep neural networks (Li et al., 2015; Nguyen et al., 2021; Mehrer et al., 2020), investigate human-machine alignment (Sucholutsky et al., 2023; Xu & Vaziri-Pashkam, 2021), and as an objective for model distillation (Tian et al., 2020; Saha et al., 2022; Zong et al., 2023). Different representational similarity measures have been proposed, including Canonical Correlation Analysis (CCA; Hotelling, 1992; Morcos et al., 2018), Singular Vector Canonical Correlation Analysis (SVCCA; Raghu et al., 2017), the previously mentioned CKA and RSA, among others (e.g., Williams et al., 2021; Ding et al., 2021; Barannikov et al., 2022; Cui et al., 2022). They can be distinguished from functional similarity measures, such as performance difference (Klabunde et al., 2025), error consistency (Geirhos et al., 2020b), and true-positive agreement (Hacohen et al., 2020), which are based on model outputs rather than internal representations (Klabunde et al., 2025), or model stitching (Bansal et al., 2021; Csiszárik et al., 2021; Merullo et al., 2023), which measures compatibility, not similarity, of internal representations (Hernandez et al., 2022).
@@
Importantly, similarity measures may differ in the invariances they evoke and the properties of representational spaces they capture (Klabunde et al., 2025). For example, local structures may be better captured by measures based on nearest neighbors (Huh et al., 2024) or localized kernels (Kornblith et al., 2019a), while capturing global similarities may be better achieved by wide kernels. Here, we make use of kernel-based CKA, as it allows probing both local and global representation structures by varying the kernel, does not require the compared representations to be of the same dimension, and is widely used (e.g., Raghu et al., 2021; Maniparambil et al., 2024; Saha et al., 2022; Zong et al., 2023; Ding et al., 2019; Baratin et al., 2021).
@@
Drivers of similarity in vision models. In contrast to language processing systems (e.g., Achiam et al., 2023; Team et al., 2023), vision models are trained with a wide range of different learning objectives. These include (a) supervised learning objectives using classification datasets such as ImageNet (Krizhevsky et al., 2012b); (b) SSL objectives with a reconstruction or multi-view matching loss (Chen et al., 2020a; Caron et al., 2021; He et al., 2022); and (c) objectives for jointly learning image and text representations (Radford et al., 2021; Zhai et al., 2023). This makes them interesting candidate models for analyzing representational similarities. The trend to larger model sizes (e.g. Dehghani et al., 2023) combined with increased dataset sizes (e.g. Schuhmann et al., 2022; Oquab et al., 2024) and multi-modality appears to be leading to a convergence of model representations (Huh et al., 2024). Similarity between otherwise identical networks trained from different random initializations increases with width (Morcos et al., 2018; Kornblith et al., 2019a), as predicted by theoretical studies of properties of neural networks in the infinite width limit (Lee et al., 2018; Matthews et al., 2018; Yang & Hu, 2021). Huh et al. (2024) hypothesize that convergence is further driven by task generality and inductive bias. Other works find that architecture and dataset (Raghu et al., 2021) play a central role in determining the representational structure of a model.
@@
Consistency of similarity. Beyond the similarity function, representational comparisons depend on both the extraction layer (Kornblith et al., 2019a; Raghu et al., 2021) and the (probe) dataset (Cui et al., 2022). For example, measuring similarity for a subset of a data distribution can yield different similarities than measuring it for the whole dataset, albeit maintaining the ranking of similarities between models (Brown et al., 2024). However, recent work demonstrated that representational similarities (e.g., measured via RSA) can be used to effectively select (downstream) task-specific models (Dwivedi & Roig, 2019; Borup et al., 2023). From a behavioral perspective, it has been shown that the ranking of the agreement of ImageNet classifier pairs generalizes to out-of-distribution data (Baek et al., 2022). Due to the limited scope of prior work (e.g., only studying a few models (Klabunde et al., 2023; Brown et al., 2024), the driving factors of the consistency of representational similarities across vision datasets have not yet been identified.
Figure 1: Pairwise similarity analysis framework. Let ${\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}}$ and ${\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}{\bm{B}}}$ be two sets of stimuli and ${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Theta}$ and ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi}$ be two sets of models. For each dataset, we use all models within the two sets to extract representations, e.g., ${\bm{Z}}^{f}_{{\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}}}$ being the features extracted from dataset ${\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}}$ using model $f$. Subsequently, for each pair of models $(f_{{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\theta}},f_{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\phi}})$, where $f_{{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\theta}}\in{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Theta}$ and $f_{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\phi}}\in{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi}$, we compute the CKA similarity between their representations (i.e., $CKA({\bm{Z}}_{{\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}}}^{f_{{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\theta}}},{\bm{Z}}_{{\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}}}^{f_{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\phi}}})$), yielding a similarity vector ${\bm{s}}_{\left({\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}},{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Theta},{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi}\right)}$. This vector can be displayed as a matrix, where each entry represents the similarity between two models for a single dataset. In the scatter plot, we contrast two such vectors computed on the same model sets but evaluated on different datasets ${\color[rgb]{0.25,0.1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.25,0.1,1}\pgfsys@color@cmyk@stroke{0.75}{0.90}{0}{0}\pgfsys@color@cmyk@fill{0.75}{0.90}{0}{0}{\bm{A}}}$ and ${\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}{\bm{B}}}$. The Pearson correlation coefficient $\rho$ between the similarities quantifies the consistency of similarities across the two datasets. The distribution of $\rho$ across all dataset pairs, ${\bm{R}}\left({\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Theta},{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi}\right)$, indicates the consistency of (relative) representational similarities across stimuli.
@@ 3 Methods
In this section, we propose a novel approach for analyzing the consistency of representational similarities across multiple datasets. Fig. 1 illustrates this framework.
@@
Computing representational similarities for a dataset. We compute the similarity between two models $f_{\theta}$ and $f_{\phi}$ for a dataset by applying CKA to their vector representations ${\bm{Z}}^{f_{\theta}}$ and ${\bm{Z}}^{f_{\phi}}$. CKA computes the similarity based on the normalized Hilbert-Schmidt Independence Criterion (Gretton et al., 2005), applied to the kernel matrices of both representations. Using CKA with a linear kernel focuses on global similarity structure, while an RBF kernel with small $\sigma$ measures local similarity structure (Kornblith et al., 2019a; Alvarez, 2023). We compared their behavior (shown in Fig. 3 and in Appx. F) and observed similar trends for both kernels. Therefore, we use CKA with a linear kernel for the remainder of this paper if not mentioned otherwise.
@@
Computing the similarity requires the full kernel matrix over the dataset, which scales quadratically with the number of data points. Hence, the exact computation becomes intractable for large datasets. However, we demonstrate in Appx. C that for CKA linear a diverse subset of $10{,}000$ samples is sufficient to accurately estimate the similarities of two models for an ImageNet-scale dataset.
@@
Representational similarities across datasets. Because image representations depend on both a model and a dataset, the representational similarities between two models may vary across different datasets. To quantify how consistent these similarities remain across datasets (i.e., the transferability of similarity relationships), we compute the correlation coefficient for pairwise model similarities (measured for a large set of model pairs) between datasets. We can then analyze these correlations for specific subsets of models grouped by their training factors.
@@
Specifically, we scrutinize how the pairwise similarities of models from two model sets, $\Theta$ and $\Phi$, correspond between two datasets, ${\bm{A}}$ and ${\bm{B}}$. By computing the pairwise representational similarities between all model pairs $(f_{\theta},f_{\phi})$, separately for each dataset, where $f_{\theta}\in\Theta$, $f_{\phi}\in\Phi$ and $f_{\theta}\neq f_{\phi}$, we obtain a similarity vector ${\bm{s}}\in\mathbb{R}^{k}$, where $k$ denotes the number of distinct model combinations. A similarity vector for two sets of models and a dataset ${\bm{X}}$ is then defined as ${\bm{s}}_{\left({\bm{X}},\Theta,\Phi\right)}=\left(\mathrm{CKA}({\bm{Z}}^{f_{\theta}}_{\bm{X}},{\bm{Z}}^{f_{\phi}}_{\bm{X}})\mid f_{\theta}\in\Theta,f_{\phi}\in\Phi,f_{\theta}\neq f_{\phi}\right)^{\top}\in\mathbb{R}^{k}$.
@@
To quantify the consistency of similarities between two datasets ${\bm{A}}$ and ${\bm{B}}$, we use the Pearson correlation coefficient between the similarity vectors ${\bm{s}}_{({\bm{A}},\Theta,\Phi)}$ and ${\bm{s}}_{({\bm{B}},\Theta,\Phi)}$, i.e., $\rho\left({\bm{s}}_{({\bm{A}},\Theta,\Phi)},{\bm{s}}_{({\bm{B}},\Theta,\Phi)}\right)$. The Pearson correlation measures the degree to which the similarity trends between models are preserved across datasets, focusing on the relative positioning of model pairs rather than the absolute similarity values. In other words, if two models $f_{\phi}$ and $f_{\theta}$ are more similar than $f_{\phi}$ and $f_{\psi}$ on one dataset, a high Pearson correlation coefficient indicates that this relationship transfers to the other dataset. Conversely, a low correlation value suggests that the models’ relationships shift significantly depending on the dataset, showing variability in processing different data distributions. Thus, the Pearson correlation coefficient is a meaningful indicator of how dataset-dependent or dataset-invariant the model similarities are.
@@
To assess the consistency of these correlations across dataset pairs, we examine the distribution of correlation coefficients for all possible dataset pairs within the set of available datasets $\mathcal{D}:{\bm{R}}(\Theta,\Phi)=\left\{\rho\left({\bm{s}}_{({\bm{A}},\Theta,\Phi)},{\bm{s}}_{({\bm{B}},\Theta,\Phi)}\right)\mid{\bm{A}},{\bm{B}}\in\mathcal{D},{\bm{A}}\neq{\bm{B}}\right\}$. The spread of the correlation coefficients across all dataset pairs indicates the consistency of the relative pairwise model similarities (similarity consistency).